OpenStudy (anonymous):

Help checking my problem. I have some idea what to do. Find the slope of the line tangent of the curve defined implicitly by x^2y + (xy)^(1/2) = 36 nicer looking equation inside.

OpenStudy (anonymous):

\[x^2y+\sqrt{xy}=36\]

OpenStudy (solomonzelman):

slope of the tangent line at x what?

OpenStudy (anonymous):

oh whoops at (2,8)

OpenStudy (solomonzelman):

so at x=2.

OpenStudy (solomonzelman):

first, you have to take the derivative (you know that) can you derive it? (Keep in mind the chain rule, multiplying times dy/dx every time you get a derivative of y)

OpenStudy (solomonzelman):

can you differentiate x^2 y first?

OpenStudy (anonymous):

that becomes: \[2xy + x^2\frac{ dy }{ dx }\]

OpenStudy (solomonzelman):

yes, correct.

OpenStudy (solomonzelman):

now the square root of \(\large\color{red}{ \sqrt{xy} }\). keep in mind the chain rule, not only for the product of x and y, but for y (the dy/dx) .

OpenStudy (solomonzelman):

Saying that. We know that \(\large\color{red}{ \frac{d}{dx} \sqrt{x} =\frac{1}{2\sqrt{x}} }\) and We know that \(\large\color{blue}{ \frac{d}{dx} xy =y+x\frac{dy}{dx} }\)

OpenStudy (solomonzelman):

So, the square root of (xy) does as follows,